OpenStudy (anonymous):

What is the solution(s) to the following system of equations? y = x2 + 4x + 9 y = x2 + 3x + 8 (4, 3) (–1, 6) (–1, 2) (4, 8)

OpenStudy (goldphenoix):

Does x2 mean \[\large x^2\]?

OpenStudy (anonymous):

Add both of the lines together, and solve for y. then plug in the x value from each of the answer until one gives the correct y value.

OpenStudy (anonymous):

@GoldPhenoix yes

OpenStudy (anonymous):

\[(x ^{2}+4x+9) + (x ^{2}+ 3x+8)\]

OpenStudy (anonymous):

is that how i add them together?

OpenStudy (anonymous):

yup and you have to add the y together too, so you'll get 2y = that. simplify and divide both sides by two.

OpenStudy (anonymous):

do i add them together by foiling?

OpenStudy (anonymous):

Foiling is only used for multiplying polynomials. You can just add the like terms.

OpenStudy (anonymous):

okay, so how would i add x^2 + x^2?

OpenStudy (anonymous):

x^2+x^2=(1+1)x^2 = 2x^2

OpenStudy (anonymous):

\[2y=2x ^{2}+7+17\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

7x but yes. Now divide by 2 and just plug in and you'll get the answer.

OpenStudy (anonymous):

\[y=x ^{2}+3.5x+\]

OpenStudy (anonymous):

whoops, plus 8.5 at the end

OpenStudy (anonymous):

\[y=x ^{2}+\frac{ 7 }{ 2 }x+\frac{ 17 }{ 2 }\]

OpenStudy (anonymous):

what you have done so far is wrong..

OpenStudy (anonymous):

if you want to solve this by substitution, you can choose to put the 1st eqn into the 2nd eqn. so wherever you see "y" in the second eqn, place the 1st eqn there..

OpenStudy (anonymous):

http://sketchtoy.com/37164441

OpenStudy (anonymous):

do you understand what i have done?

OpenStudy (anonymous):

ohh okay and then you would plug that -1 into the other equation for x?

OpenStudy (anonymous):

or either of the equations?

OpenStudy (anonymous):

yes...but you rather plug in -1 which is the value of x into either of the equations to find the value of y.

OpenStudy (anonymous):

\[-1^{2}+49-1)+9= -1^{2}+3(-1)+8\]

OpenStudy (anonymous):

oops i didnt mean 49

OpenStudy (anonymous):

i meant to put + 4(-1)

OpenStudy (anonymous):

\[-1+4(-1)+9=-1^{2}+3(-1)+8\]

OpenStudy (anonymous):

i meant you could put the value of x in either of the eqns. so \[\Large y=(-1)^2+4(-1)+9\] \[\Large y=??\]

OpenStudy (anonymous):

you get the same value when you plug it into the other eqn.

OpenStudy (anonymous):

y= 14?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so is it the second one?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw...(:

OpenStudy (anonymous):

@kausarasalley, The method I told him to do works fine.

OpenStudy (anonymous):

@JBrandonS do you get the correct answer when you use your method?

OpenStudy (anonymous):

yes.